OpenStudy (sasogeek):

if lana gives me 10 medals for good question out of 24 questions, what is the probability that she'll give me a medal for my 25th question if i ask 1 good question out of every 3 questions? please explain :) i don't know the answer but i just came up with the question which seems like a legit question too though :) help me out here guys

OpenStudy (anonymous):

So probability of asking a good question is 1/3

OpenStudy (sasogeek):

yup

OpenStudy (anonymous):

Assuming she sees this, the probability of her awarding a medal is approximately 1.00000000000 in 1.00000000000.

OpenStudy (sasogeek):

lol

OpenStudy (anonymous):

Probability that Lana gives you a medal is 10/24 = 5/12.

OpenStudy (anonymous):

That's how often she's awarded a medal in the past, but my amazing powers of foresight tell me it's definitely 1/1 for this one.

OpenStudy (anonymous):

so far it's 0/1

OpenStudy (sasogeek):

LOL but oh well, lets try to figure this out before she sees it herself lol

OpenStudy (anonymous):

lol

OpenStudy (saifoo.khan):

SPAM!

OpenStudy (anonymous):

lol

OpenStudy (sasogeek):

you wanna see what spam is? check out the medallion group

OpenStudy (saifoo.khan):

im math group mod not that group's mod. :D

OpenStudy (anonymous):

haha saif

OpenStudy (saifoo.khan):

;]

OpenStudy (anonymous):

not spamm

OpenStudy (anonymous):

Hm, I think (may be wrong) that the two events are independent to each other so just taking the product would give you a good answer.

OpenStudy (anonymous):

Legit question lol

OpenStudy (saifoo.khan):

Ok, so i delete this?

OpenStudy (sasogeek):

nope u don't :)

OpenStudy (anonymous):

Noo, that would be bad

OpenStudy (saifoo.khan):

Lol, bad things are fun to do!

OpenStudy (anonymous):

The aforementioned powers of foresight aren't powerful enough to predict the outcomes of moderator actions, so when I said 1/1....

OpenStudy (sasogeek):

medallion group = group for free medals all the time lol :P kinda fun

OpenStudy (anonymous):

Haha seriously? So everyone in Medallion Group is level 100? :-P

OpenStudy (anonymous):

NOt yet

OpenStudy (sasogeek):

i just started it lol :P you can join if u want, make a post and everyone gives u a medal :) and anyone who comments also gets a medal... kinda simple

OpenStudy (anonymous):

lol

OpenStudy (saifoo.khan):

LOL, let's report that group then! JK

OpenStudy (anonymous):

^^ meanie saifoo :(

OpenStudy (saifoo.khan):

LOL. ;)

OpenStudy (saifoo.khan):

YAY! meanies Rule!

OpenStudy (anonymous):

Haha I hunt enough undeserved medals in this group as it is ;)

OpenStudy (saifoo.khan):

That's the spirit. :D

OpenStudy (anonymous):

Saifffff awee dun do that!!(: haha u report it ill be mad at u :P

OpenStudy (anonymous):

You underserved medal hunterer you lol :P

OpenStudy (saifoo.khan):

LOL, WAIT. IM MEANIE.. THEY DO MEAN THINGS! :D

OpenStudy (anonymous):

meanieeee ^^^^ :P

OpenStudy (anonymous):

:O Saif awe u should be nice not mean!!!!!! ;)

OpenStudy (saifoo.khan):

im trying to be different.. everyone's nice everytime.. so let's be different... try being meanie! :D

OpenStudy (anonymous):

Change isnt always good

OpenStudy (anonymous):

Haha U dont want me being a meanie, u really dont saif haha ;)

OpenStudy (saifoo.khan):

WE WANT CHANGE - Obama. ;)

OpenStudy (saifoo.khan):

http://www.youtube.com/watch?v=k3rO-GP7wlE

OpenStudy (anonymous):

lolol

OpenStudy (saifoo.khan):

;]

OpenStudy (sriram):

u give 1 good question out of 3 quests hence in 24 quests u gave 8 good quests but lana smhow gives u 10 medals hence i assume she found all ur good quests good n gave u a medal for those n gave u 2 medals for 2 of the 16 bad quests n thats 2/16=1/8 hence lana will give u a medal if u gave a good quest or if she liked d bad one so i guess the probability for givin a medal is =1/3+1/8=11/24

OpenStudy (lalaly):

Lol Hi saso

OpenStudy (sasogeek):

Hi Lana :) FSM was right lol, the probability of u giving a medal for this question was 1.0 lol

OpenStudy (lalaly):

hehehe :D hes smart

OpenStudy (anonymous):

nice question saso

OpenStudy (andriod09):

@Iana